OpenStudy (anonymous):

A town has a population of 1,000 people at time t=0. For circumstances a and b, write a formula for the population, P, of the town as a function of year t. a. The population increases by 50 people a year. b. The population increases by 5% a year. **could you please explain to me how this works? I don't really get how to format the formulas :/ Thanks a bunch!! :)

OpenStudy (wolf1728):

Part a) Population = 1,000 + [(50) * years] Part b) Population = 1,000 * (1.05)^years I don't think I could explain very well how I derived those formulas. I just try to configure the formulas making changes as needed until I'm sure they work.

OpenStudy (anonymous):

so it would just be like this? a) P=1000+50t b) 1000 x (1.05)^t ?? and how come it's 1.05? thank you :) don't know if what i wrote above matches what you wrote? :/

OpenStudy (wolf1728):

Part b states that the population increases 5% per year, hence the 1.05

OpenStudy (anonymous):

ohh okay I see :) so how I rewrote what you said is correct?? a) P=1000+50t b) 1000 x (1.05)^t ? :)